Please help me solve the following
(logx)^2 - logx^2 = 3
Thanks!!!

Answer by troymuga(3)

Set u=logx
This becomes
u^2-2u=3
u^2-2u-3=0
(u-3)(u+1)=0
u=3,-1
log(x)=3
log(x)=-1
x=10^3,10^-1
x=1000,0.1